OpenStudy (anonymous):

tan^2x=sin^2x(1+tan^2x) true or false?

myininaya (myininaya):

true

myininaya (myininaya):

1+tan^2x=sec^2 sec^2x=1/cos^2x sin^2x/cos^2x=tan^2x